OpenStudy (anonymous):

Solving For X...I think I got it? Given: X/5+1=7 I did: 7+1=8, 8*5=40 which would give me X=40?

OpenStudy (cruffo):

When you move the 1 over, it will be negative: 7 - 1

OpenStudy (anonymous):

Oh.. so 7-1=6, 6*5=30, X=30?

OpenStudy (cruffo):

That looks good. You can check your answer by replacing x with 30 in the original equation: 30/5 +1 ??? 7 6 + 1 = 7 true!

OpenStudy (anonymous):

That makes alot of sense now, Thank you for the help :)

OpenStudy (cruffo):

Cheers!